OpenStudy (anonymous):

use the geometric mean find the 7th term in a geometric sequence if the 6th is term and the 8th term is 32.

OpenStudy (anonymous):

please help

OpenStudy (amistre64):

\[a_6=a_o r^5\] \[a_7=a_o r^6\] \[a_8=a_o r^7\] a geomtric mean is defined as: given 2 value, m,n, the geoMean is sqrt(nm) see if we can get the value of a7 from \(\sqrt{a_o r^5 ~a_o r^7}\) \(\sqrt{(a_o)^2 r^{5+7}}\) \(a_o\sqrt{r^{12}}\) \(a_o~r^{12/2}\) \(a_o~r^{6}\) seems to work out just fine